OpenStudy (anonymous):

Which of the following represents x squared minus sixteen all over x squared minus four x in simplified form? Answer x plus four all over x one plus four over x one minus four over x x minus four all over x

OpenStudy (anonymous):

(x^2- 16)/(x^2-4x) = (x-4)(x+4)/x(x-4) = (x+4)/x x plus four all over x